OpenStudy (anonymous):

What is the slope of a line parallel to the line with equation 2x – 5y = 9?

OpenStudy (anonymous):

slope of two parallel lines is the same.

OpenStudy (anonymous):

A 2/5 B -5/2 C -9/5 D -2/5

OpenStudy (anonymous):

first write in the form y=mx+c and find m

OpenStudy (anonymous):

it is not D

OpenStudy (anonymous):

@hoblos Is it D?

OpenStudy (anonymous):

show me how you write in the form y=mx+c

OpenStudy (hoblos):

it's not D the slopes of parallel lines are equal you have to find the slope of 2x – 5y = 9 you have to transform it to the form y = mx + b and m will be the slope

OpenStudy (hoblos):

can you do that ?

OpenStudy (anonymous):

A?

OpenStudy (hoblos):

correct